Solve by substitution or elimination
2x-3y=8
-4x+5y=-10
i got the answer (-1,-6) but I'm not sure if it's correct it would be much appreciated if you did the elimination method, thank you.

Answer by MathLover1(20849)

2x-3y=8..............1
-4x+5y=-10...........2
_________________________
2x-3y=8..............1...both sides multiply by 2
2x*2-3y*2=8*2
4x-6y=16
_____________________________________...now take 1a and 2

4x-6y=16...............1a
+
-4x+5y=-10...........2
_________________________________...add them...

4x-4x-6y+5y=16-10

-y=6....both sides multiply by -1
-y*(-1)=6*(-1)
y=-6
now find x
2x-3y=8..............1
2x-3(-6)=8
2x+18=8
2x=8-18
2x=-10
x=-5